# A conjecture on descents, inversions and the weak order on Coxeter groups

### v2

@@ 3.1. Standardization of words, inversions and descents
This is readily seen on the word $w$ as all left letters are indeed at the left-hand side of all the right letters of $w$, which are at the right-hand side of the maximal letter $n$. So $w$ does not contain an inversion $(a,c)$ with $a\in\mathcal{L}(w)$ and $c\in\mathcal{R}(w)$. The lemma follows since inversions of $u$ and $v$ are contained inversions of $w$, by definition of bipartitions. ∎
@@
A more ¨Coxeter-like’ argument to prove this lemma is with the transitivity of inversions (bipodal sets). Indeed, as $c\in\mathcal{R}(w)$, then by Definition 3.9, there is an inversion $(c,n)$ in $w$. If $(a,c)$ is also an inversion of $w$, we obtain that $(a,n)$ is an inversion by transitivity and so $a\in\mathcal{R}(w)$.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. If there is a descent in $u$ (or in $v$) between two letters $c$ and $a$, then either $a,c\in\mathcal{R}(w)$ or $a,c\in\mathcal{L}(w)$.
@@
Suppose that we have a descent in $u$ between two letters $c$ and $a$ (with $a<c$). So $u=\dots ca\dots$. Note that in particular, $u$ contains the inversion $(a,c)$. So by Lemma 3.14, if $c\in\mathcal{R}(w)$, then $a\in\mathcal{R}(w)$.
Suppose by contradiction that $c\in\mathcal{L}(w)$ and $a\in\mathcal{R}(w)$. Since $c\in\mathcal{L}(w)$, we have $u=\dots ca\dots n\dots$ by Lemma 3.14. In particular, $(a,n)\not\in\operatorname{invs}(u)$. Since $a\in\mathcal{R}(w)$, we know that $(a,n)\in\operatorname{invs}(w)$. Moreover, $\operatorname{invs}(w)=\operatorname{invs}(u)\sqcup\operatorname{invs}(v)$ by definition of bipartition, so $(a,n)\in\operatorname{invs}(v)$ and $(a,c)\notin\operatorname{invs}(v)$. By Proposition 3.5 (2) and since $a<c<n$, we must have $(c,n)\in\operatorname{invs}(v)\subset\operatorname{invs}(w)$. So $c\in\mathcal{R}(w)$ by Lemma 3.14, which is a contradiction since $c\in\mathcal{L}(w)$. ∎
@@
$\operatorname{invs}_{\mathcal{L}}(w)=\{(a,c)\in\operatorname{invs}(w)\mid a\in\mathcal{L}(w)\text{ and }b\in\mathcal{L}(w)\},$
i.e., the inversions of $w$ where both letters are left letters. Similarly, we define $\operatorname{invs}_{\mathcal{R}}(w)$ the set of right inversions of $w$, i.e., the inversions where both letters are right letters. As $\{u,v\}$ is a bipartition of $w$ and by definition of $\mathcal{L}(w)$ and $\mathcal{R}(w)$, we have:
$\operatorname{invs}_{\mathcal{L}}(w)$ $=\operatorname{invs}_{\mathcal{L}}(u)\sqcup\operatorname{invs}_{\mathcal{L}}(v),$
$\operatorname{invs}_{\mathcal{R}}(w)$ $=\operatorname{invs}_{\mathcal{R}}(u)\sqcup\operatorname{invs}_{\mathcal{R}}(v).$
Now the inversions of $u_{\mathcal{L}}$, $v_{\mathcal{L}}$, and $w_{\mathcal{L}}$ (resp. $u_{\mathcal{R}}$, $v_{\mathcal{R}}$, and $w_{\mathcal{R}}$) are just a relabeling of the left (resp. right) inversions of $u$, $v$, and $w$. This proves that $\{u_{\mathcal{L}},v_{\mathcal{L}}\}$ is a bipartition of $w_{\mathcal{L}}$ and similarly, $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ is a bipartition of $w_{\mathcal{R}}$, which proves the first statement of the proposition.
Equality (2) is readily seen: $w_{\mathcal{L}}$ is the (standardized) left part of $w$ while $w_{\mathcal{R}}$ is the (standardized) right part of $w$. If $n$ is in position $j$, all descents in positions $i<j-1$ are descents in $w_{\mathcal{L}}$ and all descents in positions $k\geq j$ are descents in $w_{\mathcal{R}}$. Moreover, there is no decent in position $j-1$ because the letter at this position is strictly smaller than $n$, the letter at position $j$.
@@
Suppose first that there is a descent in $u_{\mathcal{L}}$ that does not arise from a descent of $u$. So there is two letters $c>a$ in $\mathcal{L}(w)$ which are not consecutive in $u$, but their standardized equivalent in $u_{\mathcal{L}}$ are. So $u=\dots cd_{1}d_{2}\dots d_{k}a\dots$, where $d_{1}d_{2}\dots d_{k}$ is a non empty word on $\mathbb{N}^{*}$ constituted only of letters $d_{i}\in\mathcal{R}(w)$. By Proposition 3.17, there no descent between $c$ and $d_{1}$ and so $d_{1}>c$. Therefore $(a,d_{1})\in\operatorname{invs}(u)$ with $a\in\mathcal{L}(w)$ and $d_{1}\in\mathcal{R}(w)$, contradicting Lemma 3.14. So there is no new descent in $u_{\mathcal{L}}$.
Suppose now that there is a descent in $u_{\mathcal{R}}$ that does not arise from a descent of $u$. So there is two letters $c>b$ in $\mathcal{R}(w)$ which are not consecutive in $u$, but their standardized equivalent in $u_{\mathcal{R}}$ are. So $u=\dots ca_{1}a_{2}\dots a_{k}b\dots$, where $a_{1}a_{2}\dots a_{k}$ is a non empty word on $\mathbb{N}^{*}$ constituted only of letters $a_{i}\in\mathcal{L}(w)$. By a similar reasoning as above, we have $a_{k}<b<c$. So $(a_{k},c)\in\operatorname{invs}(u)$ with $a_{k}\in\mathcal{L}(w)$ and $c\in\mathcal{R}(w)$, which contradicts again Lemma 3.14.
In conclusion, there is no descent that are created in $u_{\mathcal{L}}$ or $u_{\mathcal{R}}$ that does not arise from the descents of $u$, hence we have $d_{R}(u)=d_{R}(u_{\mathcal{L}})+d_{R}(u_{\mathcal{R}}).$
@@
$w_{\mathcal{R}}=\operatorname{std}(6341)=4231,\,u_{\mathcal{R}}=\operatorname{std}(3461)=2341,\,v_{\mathcal{R}}=\operatorname{std}(1634)=1423.$
@@
$\tilde{w_{\mathcal{R}}}=231,\,\tilde{u_{\mathcal{R}}}=231,\,\tilde{v_{\mathcal{R}}}=123.$
We lose one descent in $w_{\mathcal{R}}$ as $4$ is the first letter and maximal by definition of the right word. In $u_{\mathcal{R}}$, the descent between $4$ and $1$ is ‘kept’ in $\tilde{u_{\mathcal{R}}}$ because the letter before $4$ is $3>1$. In $v_{\mathcal{R}}$, the descent between $4$ and $2$ is ‘lost’ because the letter before $4$ is $1<2$.
@@
As $w_{1}=n$, we know that $(a,n)\in\operatorname{invs}(w)$. Assume $u=\dots a\dots n\dots$. Then $(a,n)\not\in\operatorname{invs}(u)$. By definition of bipartitions, $\operatorname{invs}(w)=\operatorname{invs}(u)\cup\operatorname{invs}(v)$ so $(a,n)\in\operatorname{invs}(v)$ which means that $v=\dots n\dots a\dots$. The convers follows since the role of $u$ and $v$ are symmetric in the preceding argument. ∎
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in S_{n}$ with $w_{1}=n$. We prove that $d_{R}(\tilde{u})+d_{R}(\tilde{v})=d_{R}(u)+d_{R}(v)+1$. Note that this proves the proposition in general, as if $\{u,v\}$ is a bipartition of a permutation $w\in S_{n}$, then $\{u_{\mathcal{R}},v_{\mathcal{R}}\}$ is a bipartition of a certain permutation $w_{\mathcal{R}}\in S_{k}$ with $k\leq n$ such that ${w_{\mathcal{R}}}(1)=k$. When the permutation $u$ is decreased to $\tilde{u}$, all the descents of $u$ are left unchanged except the possible descent created by the letter $n$.
Suppose first that $n$ is the last letter of $u$, then $d_{R}(\tilde{u})=d_{R}(u)$. By Lemma 3.19, as all letters are the lefthand side of $n$ in $u$, they are therefore all at the right-hand side of $n$ in $v$ and $v_{1}=n$. In particular $d_{R}(\tilde{v})=d_{R}(v)-1$, which implies the desired result. Similarly, if $u_{1}=n$, then $v_{n}=n$ and we have $d_{R}(\tilde{u})=d_{R}(u)-1$ and $d_{R}(\tilde{v})=d_{R}(v)$.
@@
  • either $u=\dots bna\dots$ and $d_{R}(\tilde{u})=d_{R}(u)$; the descent is said to be kept in $\tilde{u}$.
  • Or we have $u=\dots anb\dots$ and $d_{R}(\tilde{u})=d_{R}(u)-1$, the descent is said to be lost in $\tilde{u}$.
If the descent is kept in $\tilde{u}$ and lost in $\tilde{v}$ (or the other way around), then (5) is true. We prove that the descent cannot be kept in both, nor lost in both. We proceed by contradiction.
Suppose that the descent is kept in both $\tilde{u}$ and $\tilde{v}$.
Let $a^{\prime}<b^{\prime}$ such that $v=\dots b^{\prime}na^{\prime}\dots$ (the descent is kept in $\tilde{v}$). Since $a$ is at the right-hand side of $n$ in $u$, the letter $a$ is at the left-hand side of $n$ in $v$, by Lemma 3.19. In particular $a\neq a^{\prime}$. Similarly with $b$, $a^{\prime}$, and $b^{\prime}$, we obtain that $u$ and $v$ are of the form:
@@
Note that we might have $a^{\prime}=b$ or $a=b^{\prime}$. Without loss of generality, we suppose $a^{\prime}<a$; so in particular $a^{\prime}<a<b$ (the descent is lost in $\tilde{v}$). We have $(a^{\prime},a)\in\operatorname{invs}(v)\subseteq\operatorname{invs}(w)$. Since $(a,b)\in\operatorname{invs}(u)\subseteq\operatorname{invs}(w)$, we have by transitivity of the inversion set $\operatorname{invs}(w)$ (Proposition 3.5 (1)) that $(a^{\prime},b)\in\operatorname{invs}(w)$. But $(a^{\prime},b)\notin\operatorname{invs}(u)\cup\operatorname{invs}(v)=\operatorname{invs}(w)$, which is a contradiction.
Suppose that the descent is lost in both $\tilde{u}$ and $\tilde{v}$.
@@
Without loss of generality, we suppose $a^{\prime}<a$. We obtain $a^{\prime}<a<b$. Therefore $(a^{\prime},b)\in\operatorname{invs}(u)\cap\operatorname{invs}(v)$, contradicting that $\operatorname{invs}(u)\cap\operatorname{invs}(v)=\emptyset$ ($\{u,v\}$ is bipartition of $w$). ∎
@@ 4. The case of hyperoctahedral groups (type $B$)
$[\![\bar{n},n]\!]^{*}=\{\bar{n},\dots,\bar{1}\}\cup\{1,\dots,n\},$
where $\bar{i}:=-i$. Then $W_{n}=\{\sigma\in\mathcal{S}_{[\![\bar{n},n]\!]^{*}}\mid\sigma(\bar{i})=\overline{\sigma(i)}\}$ is a subgroup of $\mathcal{S}_{[\![\bar{n},n]\!]^{*}}$, the group of permutations on $[\![\bar{n},n]\!]^{*}$. The simple reflection $\tau_{0}$ is the transposition (1¯ $(\bar{1}\ 1)\in\mathcal{S}_{[\![\bar{n},n]\!]^{*}}$ and, for $i>0$, the simple reflection $\tau_{i}$ is the product of two simple transpositions in $[\![\bar{n},n]\!]^{*}$: $\tau_{i}=(\overline{i+1}\ \bar{i})(i\ i+1)$. It is well-known that the set of reflections of $(W_{n},S)$ is:
@@
A signed permutation $\sigma\in W_{n}$, as a permutation on the set $[\![\bar{n},n]\!]^{*}$, has a one-line notation
@@
We have of course $\overline{\bar{i}}=-(-i)=i$. The symbol ‘||||’ in the notation above is used to mark the symmetry $\sigma_{\bar{i}}=\sigma(\bar{i})=\sigma(-i)=-\sigma(i)=\overline{\sigma(i)}=\overline{\sigma_{i}}$). Note that even though it is enough to keep track of only the positive positions to retrieve the whole element, it is more convenient in this article to keep track of the negative positions as well.
@@
4.1. Type $B$ permutations, inversions and standardization

### v3

@@ 3.1. Standardization of words, inversions and descents
This is readily seen on the word $w$ as all left letters are indeed at the left-hand side of all the right letters of $w$, which are at the right-hand side of the maximal letter $n$. So $w$ does not contain an inversion $(a,c)$ with $a\in{\overleftarrow{n}}(w)$ and $c\in{\overrightarrow{n}}(w)$. The lemma follows since inversions of $u$ and $v$ are contained inversions of $w$, by definition of bipartitions. ∎
@@
A more ¨Coxeter-like’ argument to prove this lemma is with the transitivity of inversions (bipodal sets). Indeed, as $c\in{\overrightarrow{n}}(w)$, then by Definition 3.11, there is an inversion $(c,n)$ in $w$. If $(a,c)$ is also an inversion of $w$, we obtain that $(a,n)$ is an inversion by transitivity and so $a\in{\overrightarrow{n}}(w)$.
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in\mathcal{S}_{n}$. If there is a descent in $u$ (or in $v$) between two letters $c$ and $a$, then either $a,c\in{\overrightarrow{n}}(w)$ or $a,c\in{\overleftarrow{n}}(w)$.
@@
Suppose that we have a descent in $u$ between two letters $c$ and $a$ (with $a<c$). So $u=\dots ca\dots$. Note that in particular, $u$ contains the inversion $(a,c)$. So by Lemma 3.16, if $c\in{\overrightarrow{n}}(w)$, then $a\in{\overrightarrow{n}}(w)$.
Suppose by contradiction that $c\in{\overleftarrow{n}}(w)$ and $a\in{\overrightarrow{n}}(w)$. Since $c\in{\overleftarrow{n}}(w)$, we have $u=\dots ca\dots n\dots$ by Lemma 3.16. In particular, $(a,n)\not\in\operatorname{Inv}(u)$. Since $a\in{\overrightarrow{n}}(w)$, we know that $(a,n)\in\operatorname{Inv}(w)$. Moreover, $\operatorname{Inv}(w)=\operatorname{Inv}(u)\sqcup\operatorname{Inv}(v)$ by definition of bipartition, so $(a,n)\in\operatorname{Inv}(v)$ and $(a,c)\notin\operatorname{Inv}(v)$. By Proposition 3.7 (2) and since $a<c<n$, we must have $(c,n)\in\operatorname{Inv}(v)\subset\operatorname{Inv}(w)$. So $c\in{\overrightarrow{n}}(w)$ by Lemma 3.16, which is a contradiction since $c\in{\overleftarrow{n}}(w)$. ∎
@@
$\operatorname{Inv}_{\overleftarrow{n}}(w)=\{(a,c)\in\operatorname{Inv}(w)\mid a\in{\overleftarrow{n}}(w)\text{ and }b\in{\overleftarrow{n}}(w)\},$
i.e., the inversions of $w$ where both letters are left letters. Similarly, we define $\operatorname{Inv}_{\overrightarrow{n}}(w)$ the set of right inversions of $w$, i.e., the inversions where both letters are right letters. As $\{u,v\}$ is a bipartition of $w$ and by definition of ${\overleftarrow{n}}(w)$ and ${\overrightarrow{n}}(w)$, we have:
$\operatorname{Inv}_{\overleftarrow{n}}(w)$ $=\operatorname{Inv}_{\overleftarrow{n}}(u)\sqcup\operatorname{Inv}_{\overleftarrow{n}}(v),$
$\operatorname{Inv}_{\overrightarrow{n}}(w)$ $=\operatorname{Inv}_{\overrightarrow{n}}(u)\sqcup\operatorname{Inv}_{\overrightarrow{n}}(v).$
Now the inversions of $u_{\overleftarrow{n}}$, $v_{\overleftarrow{n}}$, and $w_{\overleftarrow{n}}$ (resp. $u_{\overrightarrow{n}}$, $v_{\overrightarrow{n}}$, and $w_{\overrightarrow{n}}$) are just a relabeling of the left (resp. right) inversions of $u$, $v$, and $w$. This proves that $\{u_{\overleftarrow{n}},v_{\overleftarrow{n}}\}$ is a bipartition of $w_{\overleftarrow{n}}$ and similarly, $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ is a bipartition of $w_{\overrightarrow{n}}$, which proves the first statement of the proposition.
Equality (2) is readily seen: $w_{\overleftarrow{n}}$ is the (standardized) left part of $w$ while $w_{\overrightarrow{n}}$ is the (standardized) right part of $w$. If $n$ is in position $j$, all descents in positions i<j1i<j-1 are descents in $w_{\overleftarrow{n}}$ and all descents in positions $k\geq j$ are descents in $w_{\overrightarrow{n}}$. Moreover, there is no decent in position j1j-1 because the letter at this position is strictly smaller than $n$, the letter at position $j$.
@@
Suppose first that there is a descent in $u_{\overleftarrow{n}}$ that does not arise from a descent of $u$. So there is two letters $c>a$ in ${\overleftarrow{n}}(w)$ which are not consecutive in $u$, but their standardized equivalent in $u_{\overleftarrow{n}}$ are. So $u=\dots cd_{1}d_{2}\dots d_{k}a\dots$, where $d_{1}d_{2}\dots d_{k}$ is a non empty word on $\mathbb{N}^{*}$ constituted only of letters $d_{i}\in{\overrightarrow{n}}(w)$. By Proposition 3.19, there no descent between $c$ and $d_{1}$ and so $d_{1}>c$. Therefore $(a,d_{1})\in\operatorname{Inv}(u)$ with $a\in{\overleftarrow{n}}(w)$ and $d_{1}\in{\overrightarrow{n}}(w)$, contradicting Lemma 3.16. So there is no new descent in $u_{\overleftarrow{n}}$.
Suppose now that there is a descent in $u_{\overrightarrow{n}}$ that does not arise from a descent of $u$. So there is two letters $c>b$ in ${\overrightarrow{n}}(w)$ which are not consecutive in $u$, but their standardized equivalent in $u_{\overrightarrow{n}}$ are. So $u=\dots ca_{1}a_{2}\dots a_{k}b\dots$, where $a_{1}a_{2}\dots a_{k}$ is a non empty word on $\mathbb{N}^{*}$ constituted only of letters $a_{i}\in{\overleftarrow{n}}(w)$. By a similar reasoning as above, we have $a_{k}<b<c$. So $(a_{k},c)\in\operatorname{Inv}(u)$ with $a_{k}\in{\overleftarrow{n}}(w)$ and $c\in{\overrightarrow{n}}(w)$, which contradicts again Lemma 3.16.
In conclusion, there is no descent that are created in $u_{\overleftarrow{n}}$ or $u_{\overrightarrow{n}}$ that does not arise from the descents of $u$, hence we have $d_{R}(u)=d_{R}(u_{\overleftarrow{n}})+d_{R}(u_{\overrightarrow{n}}).$
@@
$w_{\overrightarrow{n}}=\operatorname{std}(6341)=4231,\,u_{\overrightarrow{n}}=\operatorname{std}(3461)=2341,\,v_{\overrightarrow{n}}=\operatorname{std}(1634)=1423.$
@@
$\widetilde{w_{\overrightarrow{n}}}=231,\,\widetilde{u_{\overrightarrow{n}}}=231,\,\widetilde{v_{\overrightarrow{n}}}=123.$
We lose one descent in $w_{\overrightarrow{n}}$ as $4$ is the first letter and maximal by definition of the right word. In $u_{\overrightarrow{n}}$, the descent between $4$ and $1$ is ‘kept’ in $\widetilde{u_{\overrightarrow{n}}}$ because the letter before $4$ is $3>1$. In $v_{\overrightarrow{n}}$, the descent between $4$ and $2$ is ‘lost’ because the letter before $4$ is $1<2$.
@@
As $w_{1}=n$, we know that $(a,n)\in\operatorname{Inv}(w)$. Assume $u=\dots a\dots n\dots$. Then $(a,n)\not\in\operatorname{Inv}(u)$. By definition of bipartitions, $\operatorname{Inv}(w)=\operatorname{Inv}(u)\cup\operatorname{Inv}(v)$ so $(a,n)\in\operatorname{Inv}(v)$ which means that $v=\dots n\dots a\dots$. The convers follows since the role of $u$ and $v$ are symmetric in the preceding argument. ∎
@@
Let $\{u,v\}$ be a bipartition of a permutation $w\in S_{n}$ with $w_{1}=n$. We prove that $d_{R}(\widetilde{u})+d_{R}(\widetilde{v})=d_{R}(u)+d_{R}(v)+1$. Note that this proves the proposition in general, as if $\{u,v\}$ is a bipartition of a permutation $w\in S_{n}$, then $\{u_{\overrightarrow{n}},v_{\overrightarrow{n}}\}$ is a bipartition of a certain permutation $w_{\overrightarrow{n}}\in S_{k}$ with $k\leq n$ such that ${w_{\overrightarrow{n}}}(1)=k$. When the permutation $u$ is decreased to $\widetilde{u}$, all the descents of $u$ are left unchanged except the possible descent created by the letter $n$.
Suppose first that $n$ is the last letter of $u$, then $d_{R}(\widetilde{u})=d_{R}(u)$. By Lemma 3.21, as all letters are the lefthand side of $n$ in $u$, they are therefore all at the right-hand side of $n$ in $v$ and $v_{1}=n$. In particular $d_{R}(\widetilde{v})=d_{R}(v)-1$, which implies the desired result. Similarly, if $u_{1}=n$, then $v_{n}=n$ and we have $d_{R}(\widetilde{u})=d_{R}(u)-1$ and $d_{R}(\widetilde{v})=d_{R}(v)$.
@@
  • either $u=\dots bna\dots$ and $d_{R}(\widetilde{u})=d_{R}(u)$; the descent is said to be kept in $\widetilde{u}$.
  • Or we have $u=\dots anb\dots$ and $d_{R}(\widetilde{u})=d_{R}(u)-1$, the descent is said to be lost in $\widetilde{u}$.
If the descent is kept in $\widetilde{u}$ and lost in $\widetilde{v}$ (or the other way around), then (5) is true. We prove that the descent cannot be kept in both, nor lost in both. We proceed by contradiction.
Suppose that the descent is kept in both $\widetilde{u}$ and $\widetilde{v}$.
Let $a^{\prime}<b^{\prime}$ such that $v=\dots b^{\prime}na^{\prime}\dots$ (the descent is kept in $\widetilde{v}$). Since $a$ is at the right-hand side of $n$ in $u$, the letter $a$ is at the left-hand side of $n$ in $v$, by Lemma 3.21. In particular $a\neq a^{\prime}$. Similarly with $b$, $a^{\prime}$, and $b^{\prime}$, we obtain that $u$ and $v$ are of the form:
@@
Note that we might have $a^{\prime}=b$ or $a=b^{\prime}$. Without loss of generality, we suppose $a^{\prime}<a$; so in particular $a^{\prime}<a<b$ (the descent is lost in $\widetilde{v}$). We have $(a^{\prime},a)\in\operatorname{Inv}(v)\subseteq\operatorname{Inv}(w)$. Since $(a,b)\in\operatorname{Inv}(u)\subseteq\operatorname{Inv}(w)$, we have by transitivity of the inversion set $\operatorname{Inv}(w)$ (Proposition 3.7 (1)) that $(a^{\prime},b)\in\operatorname{Inv}(w)$. But $(a^{\prime},b)\notin\operatorname{Inv}(u)\cup\operatorname{Inv}(v)=\operatorname{Inv}(w)$, which is a contradiction.
Suppose that the descent is lost in both $\widetilde{u}$ and $\widetilde{v}$.
@@
Without loss of generality, we suppose $a^{\prime}<a$. We obtain $a^{\prime}<a<b$. Therefore $(a^{\prime},b)\in\operatorname{Inv}(u)\cap\operatorname{Inv}(v)$, contradicting that $\operatorname{Inv}(u)\cap\operatorname{Inv}(v)=\emptyset$ ($\{u,v\}$ is bipartition of $w$). ∎
@@ 4. The case of hyperoctahedral groups (type $B$)
$\bar{n},n]\!]^{*}=\{\bar{n},\dots,\bar{1}\}\cup\{1,\dots,n\},$
[…]
@@
A signed permutation $\sigma\in W_{n}$, as a permutation on the set $\bar{n},n]\!]^{*}$, has a one-line notation
@@
We have of course $\overline{\bar{i}}=-(-i)=i$. The symbol ‘$|$’ in the notation above is used to mark the symmetry $\sigma_{\bar{i}}=\sigma(\bar{i})=\sigma(-i)=-\sigma(i)=\overline{\sigma(i)}=\overline{\sigma_{i}}$). Note that even though it is enough to keep track of only the positive positions to retrieve the whole element, it is more convenient in this article to keep track of the negative positions as well.
@@
4.1. Type $B$-permutations, inversions and standardization